OpenStudy (anonymous):

A region R is enclosed by the coordinate axes and the graph of y=k(x-5)^2, k>0. When this region is revolved around the x-axis, the solid formed has a volume of 2500pi cubic units. What is the value of k? Answer: k=2

OpenStudy (zehanz):

To get started, here is a graph of the situation (I took k=0.1, but it can be any positive value)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

do i need to use integrals?

OpenStudy (zehanz):

Yes, you do. Do you know the integral for a solid of revolution (about the x-axis)?

OpenStudy (anonymous):

no:/

OpenStudy (zehanz):

If you have a function f(x) this formula is \(V=\pi\int_{a}^{b}(f(x)^2 dx\)

OpenStudy (zehanz):

I think you need to know this formula to be able to solve the problem. Plug in your function: \(2500\pi=\int_{0}^{5}(k(x-5)^2)^2dx\)

OpenStudy (zehanz):

Oops, forgot the pi on the RHS! It cancels, anyway :D

OpenStudy (zehanz):

Could you try to do some work on the integrand before integrating?

OpenStudy (anonymous):

foil the (x-5)^2?

OpenStudy (zehanz):

That would give you a lot of extra work! Just keep the brackets, x-5 only means that the graph has shifted 5 units to the right, and has no other meaning. You can keep it while integrating, but first you will have to square k and (x-5)²...

OpenStudy (zehanz):

So we now have: \(2500\pi=\pi\int_{0}^{5}k^2(x-5)^4dx\)

OpenStudy (zehanz):

Simplify by canceling the pi and putting k² before the integral sign (it is a constant): \(2500=k^2\int_{0}^{5}(x-5)^4dx\)

OpenStudy (anonymous):

okay

OpenStudy (zehanz):

So now you have to integrate \((x-5)^4\). Do you know how to integrate \(x^n\) ?

OpenStudy (anonymous):

i am confused about what to do with the exponent

OpenStudy (zehanz):

Use the rule \(\int x^ndx=\dfrac{1}{n+1}x^{n+1}\)

OpenStudy (zehanz):

Just remember to use x-5 instead of x, everything is shifted 5 units to the right.

OpenStudy (anonymous):

what about the k^2?

OpenStudy (zehanz):

Leave it for now, we'll handle it later.

OpenStudy (anonymous):

so (x-5)^4 will be (x^2/2 - 5x)^4?

OpenStudy (zehanz):

No, it will be \(\dfrac{1}{5}(x-5)^5\)

OpenStudy (zehanz):

Remember the rule and set n = 4.

OpenStudy (anonymous):

whats dfrac?

OpenStudy (zehanz):

Don't you see a nice fraction? It is supposed to be like this:

OpenStudy (anonymous):

ooh okay got it

OpenStudy (anonymous):

now what do i do?

OpenStudy (zehanz):

Solve this:

OpenStudy (zehanz):

It will give you k=2. Hope you'll succeed. Sorry, have to go now...

OpenStudy (anonymous):

got it thanks a lot!:)